# VMRNN: Integrating Vision Mamba and LSTM for Efficient and Accurate Spatiotemporal Forecasting

### v2

@@ Abstract
Combining CNNs or ViTs, with RNNs for spatiotemporal forecasting, has yielded unparalleled results in predicting temporal and spatial dynamics. However, modeling extensive global information remains a formidable challenge; CNNs are limited by their narrow receptive fields, and ViTs struggle with the intensive computational demands of their attention mechanisms. The emergence of recent Mamba-based architectures has been met with enthusiasm for their exceptional long-sequence modeling capabilities, surpassing established vision models in efficiency and accuracy, which motivates us to develop an innovative architecture tailored for spatiotemporal forecasting. In this paper, we propose the VMRNN cell, a new recurrent unit that integrates the strengths of Vision Mamba blocks with LSTM. We construct a network centered on VMRNN cells to tackle spatiotemporal prediction tasks effectively. Our extensive evaluations show that our proposed approach secures competitive results on a variety of tasks while maintaining a smaller model size. Our code is available at https://github.com/yyyujintang/VMRNN-PyTorch.
@@ 1 Introduction
Recently, Structured State Space Models (SSMs) [12, 13] have emerged as notably efficient and effective in modeling extensive sequences. Mamba [9] positions itself as a potential breakthrough for addressing long-range dependencies in various tasks, which innovatively introduces the selectivity of the input sequence and uses the scan method. Unlike transformers, which often exhibit quadratic scaling for sequence length, Mamba maintains a linear or near-linear scaling, all the while adeptly handling long-range dependencies. This attribute has catapulted them to the forefront of continuous long-sequence data analysis, achieving state-of-the-art results in fields such as natural language processing and genomic analysis. A series of pioneering studies [34, 30, 58, 65, 16, 29, 60, 8, 15] have begun to investigate the utility of Mamba models within the vision sector, showcasing their versatility across a broad spectrum of vision-based applications. These investigations into Mamba’s application—ranging from basic image recognition to complex segmentation tasks—have yielded encouraging outcomes, underscoring the significant potential that SSMs hold in revolutionizing vision tasks.
@@ 3.1 Overall Architecture
For the VMRNN-B model, the VMRNN layer processes the embedded image patches, along with the previous time step’s hidden state and cell state , to generate the current hidden state and cell state . As illustrated in Fig. 2(a), is replicated, producing two versions: one is directed to the reconstruction layer, and the other, in conjunction with , serves the VMRNN layer for the subsequent time step. For VMRNN-B, the architecture primarily relies on the stacking of VMRNN layers. For the VMRNN-D variant, we incorporate more VMRNN Cells and introduce Patch Merging and Patch Expanding layers, as outlined in [2]. The Patch Merging layer is employed for downsampling, effectively reducing the spatial dimensions of the data, which aids in reducing computational complexity and capturing more abstract, global features. Conversely, the Patch Expanding layer is used for upsampling, which increases the spatial dimensions, facilitating the restoration of detail and enabling precise localization of features in the reconstruction phase. The notations and denote the hidden and cell states at time step from layer , respectively. Ultimately, the reconstruction layer takes the hidden state from the VMRNN layer and scales it back to the input size, generating the predicted frame for the next time step.
@@ 3.2 VMRNN Module
ConvLSTM [41] incorporates convolutional operators into the information transition from both input and state and state to state, effectively addressing the limitations of FC-LSTM [44] in handling spatio-temporal data.
VMRNN Module removes all weights and biases in ConvLSTM to obtain Eqn. 3:
In Eqn. 3, . Then these three gates: , , are fused into one filter gate .
We propose the VMRNN Module, which is detailedly presented in Fig. 3 (a). In VMRNN, the long-term and short-term temporal dependencies are captured by updating the information of cell states , and hidden states is updated from a horizontal perspective. And the VSS Blocks vertically learn spatial dependencies.
We show the key equations of VMRNN in Eqn. 6, where VSB means the VSS Blocks in Sec. 3.3 and LP is short for the Linear Projection:
@@ 3.3 VSS Block
VSS block addresses the challenges associated with 2D image data by employing 2D-selective-scan (SS2D), as illustrated in Fig. 2 (c). This approach unfolds image patches in four distinct directions: from the top-left to the bottom-right, from the bottom-right to the top-left, from the top-right to the bottom-left, and from the bottom-left to the top-right, creating four distinct sequences, as depicted in the Scan Expand Stage. Subsequently, each feature sequence(scan) will be processed through the S6 Block. Finally, these sequences are reconfigured back into individual images, as depicted in the Scan Merge Stage. Given input feature , the output feature of SS2D can be written as:
@@ 4.2 Main results
Moving MNIST We present the quantitative outcomes in Table 2, where our VMRNN model demonstrates notably superior performance compared to all other evaluated models. We report the results of previous research directly. On the Moving MNIST dataset, VMRNN not only achieves obviously lower MSE but also secures higher SSIM scores, significantly surpassing earlier methods by a substantial margin and archives 6.8% improvement over SwinLSTM according to MSE.
@@ 4.3 Ablation Study
The Convolution Layer. The role of the VSS Block convolution layer is to decode the spatiotemporal representations extracted by the VMRNN cell. We conduct experiments on depth-wise convolutions (DW Conv), convolution, and depth-wise convolutions with dilations (DW-D Conv) to analyze the impact of different decoding methods. Following [46], we combine DW Conv-DW-D Conv-1×1 Conv to model the large kernel convolutions, Table 6 shows that the depth-wise convolution performs much better than the other two methods.
@@ 4.4 Visualization
We present qualitative results of VMRNN on Moving MNIST in Fig. 6, TaxiBJ in Fig. 7, and KTH in Fig. 4. For all datasets, the first line is the input, the second line is the ground truth, and the third line is the prediction of VMRNN. For TaxiBJ, we add the fourth line to show the difference between prediction and target. The visualization results reveal that our VMRNN model delivers impressive predictive performance, maintaining high accuracy even across extended prediction horizons.
@@ 5 Conclusion
In this work, we introduce VMRNN, a novel approach that integrates LSTM architecture with VSS Blocks to tackle spatiotemporal forecasting challenges. Through rigorous evaluation across diverse datasets, VMRNN has proven its prowess by delivering superior performance while maintaining a smaller model size. This advancement is attributed to the model’s enhanced capability to learn and leverage global spatial dependencies with linear complexity, enabling a more refined understanding of spatiotemporal dynamics. The findings of this research affirm the potential of VMRNN as a leading-edge method for spatiotemporal prediction, setting a new strong baseline for future explorations in the field.

### v3

@@ Abstract
Combining Convolutional Neural Networks (CNNs) or Vision Transformers(ViTs) with Recurrent Neural Networks (RNNs) for spatiotemporal forecasting has yielded unparalleled results in predicting temporal and spatial dynamics. However, modeling extensive global information remains a formidable challenge; CNNs are limited by their narrow receptive fields, and ViTs struggle with the intensive computational demands of their attention mechanisms. The emergence of recent Mamba-based architectures has been met with enthusiasm for their exceptional long-sequence modeling capabilities, surpassing established vision models in efficiency and accuracy, which motivates us to develop an innovative architecture tailored for spatiotemporal forecasting. In this paper, we propose the VMRNN cell, a new recurrent unit that integrates the strengths of Vision Mamba blocks with LSTM. We construct a network centered on VMRNN cells to tackle spatiotemporal prediction tasks effectively. Our extensive evaluations show that our proposed approach secures competitive results on a variety of tasks while maintaining a smaller model size. Our code is available at https://github.com/yyyujintang/VMRNN-PyTorch.
@@ 1 Introduction
Recently, Structured State Space Models (S4) [12, 13] have emerged as notably efficient and effective in modeling extensive sequences. Mamba [9] positions itself as a potential breakthrough for addressing long-range dependencies in various tasks, which innovatively introduces the selectivity of the input sequence and uses the scan method. Unlike transformers, which often exhibit quadratic scaling for sequence length, Mamba maintains a linear or near-linear scaling, all the while adeptly handling long-range dependencies. This attribute has catapulted them to the forefront of continuous long-sequence data analysis, achieving state-of-the-art results in fields such as natural language processing and genomic analysis. A series of pioneering studies [34, 30, 58, 65, 16, 29, 60, 8, 15] have begun to investigate the utility of Mamba models within the vision sector, showcasing their versatility across a broad spectrum of vision-based applications. These investigations into Mamba’s application—ranging from basic image recognition to complex segmentation tasks—have yielded encouraging outcomes, underscoring the significant potential that SSMs hold in revolutionizing vision tasks.
@@ 3.1 Overall Architecture
For the VMRNN-B model, the VMRNN layer processes the embedded image patches, along with the previous time step’s hidden state and cell state , to generate the current hidden state and cell state . As illustrated in Fig. 2(a), is replicated, producing two versions: one is directed to the reconstruction layer, and the other, in conjunction with , serves the VMRNN layer for the subsequent time step. For VMRNN-B, the architecture primarily relies on the stacking of VMRNN layers. For the VMRNN-D variant, we incorporate more VMRNN Cells and introduce Patch Merging and Patch Expanding layers, as outlined in [2]. The Patch Merging layer is employed for downsampling, effectively reducing the spatial dimensions of the data, which aids in reducing computational complexity and capturing more abstract, global features. Conversely, the Patch Expanding layer is used for upsampling, which increases the spatial dimensions, facilitating the restoration of detail and enabling precise localization of features in the reconstruction phase. Ultimately, the reconstruction layer takes the hidden state from the VMRNN layer and scales it back to the input size, generating the predicted frame for the next time step.
@@ 3.2 VMRNN Module
We propose the VMRNN Module, detailed in Fig. 2 (a). In VMRNN, the long-term and short-term temporal dependencies are captured by updating the information of cell states , and hidden states are updated from a horizontal perspective. And the VSS Blocks vertically learn spatial dependencies. We show the key equations of VMRNN in Eqn. 6, where VSB means the VSS Blocks in Sec. 3.3 and LP is short for the Linear Projection:
@@ 3.3 VSS Block
VSS block first recovers linear projections to the image shape. (From [B, L, C] to [B, H, W, C]). Then VSS block addresses the challenges associated with 2D image data by employing 2D-selective-scan (SS2D), as illustrated in Fig. 2 (c). This approach unfolds image patches in four distinct directions: from the top-left to the bottom-right, from the bottom-right to the top-left, from the top-right to the bottom-left, and from the bottom-left to the top-right, creating four distinct sequences, as depicted in the Scan Expand Stage. Subsequently, each feature sequence(scan) will be processed through the S6 Block. Finally, these sequences are reconfigured back into individual images, as depicted in the Scan Merge Stage. Given input feature , the output feature of SS2D can be written as:
@@ 4.2 Main results
Moving MNIST We present the quantitative outcomes in Table 2, where our VMRNN model demonstrates notably superior performance compared to all other evaluated models. We report the results of previous research directly. On the Moving MNIST dataset, VMRNN not only achieves obviously lower MSE but also secures higher SSIM scores, significantly surpassing earlier methods by a substantial margin and archives 6.8% improvement over SwinLSTM.
@@ 4.3 Ablation Study
The Convolution Layer. The role of the VSS Block convolution layer is to decode the spatiotemporal representations extracted by the VMRNN cell. We conduct experiments on depth-wise convolutions (DW Conv), convolution, and depth-wise convolutions with dilations (DW-D Conv) to analyze the impact of different decoding methods. Following [46], we combine DW Conv-DW-D Conv-1×1 Conv to model the large kernel convolutions, Table 6 shows that DW Conv performs much better than the other two methods.
@@ 4.4 Visualization
We present qualitative results of VMRNN on Moving MNIST in Fig. 6, TaxiBJ in Fig. 7, and KTH in Fig. 4. For all datasets, the first line is the input, the second line is the ground truth, and the third line is the prediction of VMRNN. For TaxiBJ, we add the fourth line to show the difference between prediction and target. The visualization results reveal that VMRNN delivers impressive predictive performance, maintaining high accuracy across extended horizons.
@@ 5 Conclusion
In this work, we introduce VMRNN, a novel approach that integrates LSTM architecture with VSS Blocks to tackle spatiotemporal forecasting challenges. Through rigorous evaluation across diverse datasets, VMRNN has proven its prowess by delivering superior performance while maintaining a smaller model size. This advancement is attributed to the model’s enhanced capability to learn and leverage global spatial dependencies with linear complexity, enabling a more refined understanding of spatiotemporal dynamics. The findings show that VMRNN sets a new strong baseline for future explorations in the field.
6 Acknowledgements
This work was supported by the National Natural Science Foundation of China (No. 62306257). The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of the National Natural Science Foundation.